A 100-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 20 ft longer that the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace

Answer by mananth(16949)

A 100-ft diagonal brace on a bridge connects a support of the center of the bridge to a side support on the bridge. The horizontal distance that it spans is 20 ft longer that the height that it reaches on the side of the bridge. Find the horizontal and vertical distances spanned by this brace
...
The diagonal (d) horizontal span (h)and the vertical support(v) form a right triangle
...
h=(v+20)
d^2=v^2+h^2
(100)^2=v^2+(v+20)^2
...
100^2=v^2+v^2+40v+400
2v^2+40v+400-10000=0
2v^2+40v-9600=0
/2
v^2+20v-4800=0
v^2+80v-60x-4800=0
v(v+80)-60(v+80)=0
(v+80)(v-60)=0
v = 60 feet the vertical height
horizontal span = 80 feet
...
m.ananth@hotmail.ca